A lorry is loaded with two different types of boxes. If 150 of box A and 115
of box B are loaded onto the truck, its capacity of 10 tonnes is reached. If
300 of box A are loaded , then the truck can only take 30 of box B before
the capacity of 10 tonnes is reached. Find the weight of each box.

Answer by macston(5194)

.
150A+115B=10 tonnes . Multiply this by 2.
300A+230B=20 tonnes
.
300A+30B=10 tonnes , Subtract this from above
_____________________
200B=10 tonnes
B=10tonnes/200
B=1/20 tonne
.
150A+115B=10 tonnes
150A+115(1/20 tonne)=10 tonnes
150A+115/20 tonnes=10 tonnes
150A=200/20 tonnes-115/20 tonnes
150A=85/20 tonnes
A=85/3000 tonnes
A=17/600 tonnes
.
ANSWER: Box A weighs 17/600 tonne; Box B weighs 1/20 tonne
.
CHECK:
300A+30B=10 tonnes
300(17/600)+30(1/20)=10 tonnes
(17/2)+(3/2)=10 tonnes
20/2 tonnes=10 tonnes
10 tonnes=10 tonnes
.